OpenStudy (anonymous):

Da Eum is 5 years younger than her sister. How old will the both be when: A. The sum of their age is at most 57 years B. Da Eum sister is at least 11/10 as old as Da Eum

OpenStudy (anonymous):

the choices

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

oh, it no choices, you suppose to solve it

OpenStudy (anonymous):

ok hol up o.O

OpenStudy (anonymous):

you need to make an equation so lets say Da Eum is x yeas old and her sister is x-5

OpenStudy (anonymous):

therefore x+(x-5)=57 would be used to solve option A

OpenStudy (anonymous):

I don't get it, isn't it \[x-5\ge57\]

OpenStudy (anonymous):

it cannot just be only x-5 because that is the age of Da Eum younger sister

OpenStudy (anonymous):

Da Eum has to be x itself with her sister 5 years younger (x-5)

OpenStudy (anonymous):

and the question says at most 57 meaning it cannot be >57

OpenStudy (anonymous):

\[\le \]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

So it x+(x-5)\[\le 57\]

OpenStudy (anonymous):

yeah that is better\[x+(x-5)\le57\]

OpenStudy (anonymous):

so solve for x to find individual ages